OpenStudy (anonymous):

think for this one ...If money is invested for a short period of time, simple interest is paid using the following formula. I = Prt P represents the money invested, or principal amount, r is the simple interest rate, and t represents the time, in years, the money is invested. Given P = $3,000, r = 4%, and I = $840, how long was the money invested?

OpenStudy (anonymous):

t=I/(Pr) t=840/(3000*.04)

OpenStudy (anonymous):

7 years

OpenStudy (anonymous):

do you see how I did that?